How to solve 3x^2-27/x^2+3x
Thank you!

Answer by Jk22(389)

factorization to eventually simplify : 3x^2-27=3(x^2-9)=3(x+3)(x-3)
and x^2+3x=x(x+3), hence this fraction is simplifiable by x+3
:

3x^2-27/x^2+3x = 3(x-3)/x = 3(1-3/x)